# Immersed figure-8 annuli and anyons

### v2

@@ 2.3 Superselection sectors and fusion spaces
Information convex sets detect superselection sectors and fusion spaces. In 2D, an embedded annulus $X$ detects the anyon types ${\mathcal{C}}=\{1,a,b,\cdots\}$, and here $1$ denotes the vacuum sector. The 2-hole disk (Fig. 2 (a)) detects the fusion spaces. These follow from the structure theorems: simplex theorem and Hilbert space theorem in the original paper [14]. The simplex theorem states that the information convex set of the embedded annulus, $\Sigma(X)$, is a simplex:
@@
More generally, the simplex theorem works for any sectorizable region [15]. Sectorizable is a topological condition that says $S$ contains disjoint subsets $S_{1}$ and $S_{2}$, each of which can deform back to $S$ by a sequence of elementary extensions. In 2D, immersed annuli and disks are the only sectorizable regions. A Möbius strip embedded in $RP^{2}$ is not sectorizable, even though it is the thickening of some loop.
@@ 2.4 Quantum dimensions
Let $Z$ be an immersed annulus. Let $Z=BCD$, where $BD=\partial Z$ is the thickened boundary and $C=Z\setminus\partial Z$ is the interior; $B$ and $D$ each is a union of two disks, as illustrated in Fig. 3. For any extreme point $\rho^{h}_{Z}\in\Sigma(Z)$,
@@
Figure 3: The partition of an immersed annulus $Z=BCD$ used in defining quantum dimension (Definition 1). We first map the immersed annulus to a topological annulus by some homeomorphism indicated by the arrow. We then partition the topological annulus as illustrated.
@@
According to Lemma 4, any one of the conditions 1, 2, 3, and 4 can be used to determine if a state in $\Sigma(Z)$ is an extreme point associated with Abelian superselection sector.
@@
First, suppose that $\rho_{Z}$ is an extreme point of $\Sigma(Z)$. According to the definition of quantum dimension, $\Delta(B,C,D)_{\rho}=0$ for the partition in Fig. 3. It follows from Lemma 12 in Appendix A that any one of conditions 1, 2, 3, 4 implies the rest of them. (Lemma 12 in the appendix is a more general statement about the equivalence of definitions for quantum dimension, works beyond 2D setups.) Second, none of 1, 2, 3, 4 is true if $\rho_{Z}=\sum_{h}p_{h}\,\rho^{h}_{Z}$ is a non-extreme point. This is because a non-extreme point has contribution $\sum_{h}p_{h}\ln\left(\frac{1}{p_{h}}\right)>0$ added to $\Delta$s and $I$ in conditions 1, 2, 3, 4. This contribution is always positive for non-extreme points. These two considerations derive Lemma 4. ∎
@@ 3 Fun with figure-8
We emphasize that the conjecture is made in the context where the reference state is defined on sphere $\mathbf{S}^{2}$. An alternative version of the conjecture requires the reference state $\sigma$ defined on a disk. These two versions are equivalent due to the ability to “complete” the disk into a sphere. One may wonder if it is possible to relax the support of the reference state further. This is impossible. If the reference state is defined on an annulus, there may not be an Abelian sector, even for an embedded annulus; this is due to topological defects, see Section 5.1. We shall prove Conjecture 5 for Abelian anyon theory in Section 4.
@@
The weaker statement $|{\mathcal{C}}_{8}|=|{\mathcal{C}}|$, i.e., the number of extreme points of $\Sigma(X_{8})$ equals the number of anyon species is true; we will discuss this in Section 5. If the two information convex sets $\Sigma(X_{8})$ and $\Sigma(X)$ are isomorphic in the most natural manner, we expect an invertible map between them that preserves the quantum dimensions. Then, in addition to $|{\mathcal{C}}|=|{\mathcal{C}}_{8}|$, we should have an Abelian sector of ${\mathcal{C}}_{8}$ that is mapped to the vacuum $1\in{\mathcal{C}}$. This is one way to motivate Conjecture 5. (It is a further question whether there is a “canonical vacuum” on ${\mathcal{C}}_{8}$.)
@@ 3.1 Tunneling trick
We introduce a tunneling trick. The essential observation is that tunneling a figure-8 through a strip can add a twist to the strip. If the “figure-8” region carries an Abelian sector, this process generates an isomorphism between information convex sets associated with homeomorphic topological regions in possibly distinct regular homotopy classes. We first explain the tunneling process at a topological level. After that, we discuss the details of entanglement bootstrap and explain what tunneling does on the quantum state (density matrices) and why an Abelian sector in ${\mathcal{C}}_{8}$ plays a role.
@@
What does the tunneling process in Fig. 5 mean in entanglement bootstrap? It represents a transformation of the quantum state (density matrix) on the region containing the strip. Let the region containing the original strip be $A$ and the region containing the strip after tunneling be $A^{\prime}$, (see Fig. 6). Suppose, the state on “figure-8” $C$ is an extreme point $\rho^{\mu}_{C}$, with $\mu\in{\mathcal{C}}_{8}$. The transformation, as we shall discuss, must be a quantum channel that depends on $\mu$:
@@
Figure 6: Analyzing the attachment (detachment) of an immersed figure-8 annulus $C$ ($C^{\prime}$). Only part of $A$, $A^{\prime}$ are shown. (a) A partition for merging. (b) and (c) have special usage when the state on “figure-8” is an Abelian extreme point.
@@
The derivation of Eq. (16) is as follows. We first extend the state $\alpha_{ABC}\in\Sigma_{[\rho^{\mu}_{C}]}(ABC)$ to a state in $\Sigma(ABCD)$, as in Fig. 6(b). If $\mu$ is Abelian, by Lemma 4, we have $\Delta(B,C,D)_{\rho^{\mu}}=0$. Then by strong subadditivity, $I(A:C|B)_{\alpha}\leq\Delta(B,C,D)_{\rho^{\mu}}=0$. Because of its quantum Markov state structure, $\alpha_{ABC}$ can be recovered from $\alpha_{AB}$ by a quantum channel independent of $\alpha_{AB}$ (but depend on $\rho^{\mu}$). The fact that this $\alpha$-independent quantum channel reverses the partial trace ${\rm Tr}_{C}$ implies that $\Gamma^{\rm att}$ is an isomorphism. For the same reason, the detachment map $\Gamma^{\rm det}$ is an isomorphism too. (The state exists by merging. The reversibility follows from $I(A^{\prime}:C^{\prime}|B^{\prime})=0$.)
@@
Here, a 0-handle is a disk (the round part of the figure), and a 1-handle is the strip (thick lines in the figure). Both 0-handles and 1-handles are topological balls. The difference is the way they attach. This way of illustration will be convenient in the following.
@@ 3.2 Abelian sector on “figure-8” implies 2D strong isomorphism
Objects participating in the boundary connected sum can slide around each other since they are connected by 1-handles. (Such sliding works for both embedded regions and the more general classes of immersed regions.) It follows that we can smoothly deform $\Omega$ on $\mathbf{S}^{2}$ such that it looks like the left figure of Fig. 7. It is a 0-handle with $12k+m-$ 1 1-handles attached in the order (around the 0-handle) indicated in the figure. Some of the 1-handles have twists. Upon the removal of possible twists, $\Omega$ becomes $\Omega^{\star}$ in the right figure of Fig. 7. In $\Omega^{\star}$, the 1-handles form two groups: $(m-1)$ of them are separated from others (dark brown), and there are $k$ pairs of 1-handles (dark blue), the two 1-handles in each pair “cross” each other. Importantly, in both figures in Fig. 7, we can arrange the 1-handles in the same order, circling around the 0-handle (the chunky gray disk), thanks to the ability to slide the components around each other.
@@ 4 Proofs in the context of 2D Abelian anyon theory
We consider a sequence of operations applying to a state on “figure-8” $X_{8}$ carrying an arbitrary $\mu\in{\mathcal{C}}_{8}$; see Fig. 8. We denote this extreme point of $\Sigma(X_{8})$ as $\rho^{\mu}_{X_{8}}$. We draw $X_{8}$ in such a way that part of the “figure-8” annulus is blown up. The region is like a “watch with a twisted belt”. We shall explain the steps in Fig. 8 and the fact that they imply a useful condition Eq. (23) on the $\rho^{\mu}_{X_{8}}$. Below are the details.
@@
The key ingredient of the above proof is that transporting an anyon along the figure-8 does not change its type; here, we assume that the figure-8 region is immersed in a ball or a sphere. We emphasize again that the crucial input is the anyon transportation property in the neighborhood of the immersed region, and it is not essential to assume the immersion is in a ball or sphere if we know this property. This observation will be useful later in Section 5.
@@ 5 Transportation experiment
We discuss a thought experiment, “transportation experiment,” which reveals the physical connection between immersed annuli with topological defects. The entanglement bootstrap setup in this section is relaxed compared with that in Sections 3 and 4. Let us consider 2D for simplicity. We allow the reference state to have “defect points” where the axioms may be violated. (Alternatively, we consider a reference state defined on a manifold $N$ with nontrivial topology, e.g., a $k$-hole disk.) The setup is illustrated in Fig. 9.
@@
We want to transport an anyon along a closed loop ${\mathcal{L}}$ and back. In entanglement bootstrap, this is done by transporting a “test annulus” $X$, which is a sufficiently small embedded annulus that detects the anyon type by the information convex set $\Sigma(X)$. If we transport the test annulus $X$ along closed loop ${\mathcal{L}}$ and back to its original position, we have an automorphism of the information convex set determined by the loop ${\mathcal{L}}$: [^11] Accurately speaking, the automorphism is determined by a loop ${\mathcal{L}}$ with a chosen orientation. We may denote an oriented loop as $\vec{\mathcal{L}}$. However, we shall use ${\mathcal{L}}$ for simplicity. Quantities of interest below, such as the subset of anyons not permuted by the transportation, do not depend on the orientation. (For a discussion of automorphisms of information convex sets in broader setups, see Appendix B.)
@@
In general, the automorphism $\Phi({\mathcal{L}})$ can permute extreme points of $\Sigma(X)$, and thus, it permutes anyon types. In other words, for each ${\mathcal{L}}$, we have an automorphism of labels $\varphi_{\mathcal{L}}:{\mathcal{C}}\to{\mathcal{C}}$. The automorphism depends only on the topological class of ${\mathcal{L}}$, i.e., smoothly deforming ${\mathcal{L}}$ without passing the defect points preserves $\Phi$ and $\varphi$. Explicit examples come from topological defects, which we recall in Section 5.1. Throughout this section, we consider loops ${\mathcal{L}}$ whose thickening ($X_{\mathcal{L}}$) is an immersed annulus. [^12] This excludes some loops on nonorientable manifolds. For instance, the noncontractible loop on $RP^{2}$ can be thickened to an Möbius strip but not an annulus. We define
@@ 5.1 Topological defects: an example
A well-known example of topological defect is the one that permutes $e$ and $m$ in the toric code model [34]. The anyon types of the toric code are ${\mathcal{C}}=\{1,e,m,f\}$. Their quantum dimensions are $d_{1}=d_{e}=d_{m}=d_{f}=1$. Topological defects in this model are non-Abelian ${\mathcal{C}}^{\prime}=\{\sigma_{+},\sigma_{-}\}$ with $d_{\sigma_{+}}=d_{\sigma_{-}}=\sqrt{2}$. Anyons $e$ and $m$ are permuted when they are transported around the defect (either $\sigma_{+}$ or $\sigma_{-}$). While defects are most familiar in the context where global symmetries are present [35], it is understood that the phenomenon of permuting anyons does not go away under arbitrary local perturbations, not preserving any symmetry.
@@
The presence of a defect affects information convex sets. This is illustrated in Fig. 10 for the toric code example. The annulus $X^{\prime}$ surrounds the defect, and its information convex set $\Sigma(X^{\prime})$ has two extreme points. The quantum dimension for each extreme point, defined according to Definition 1, agrees with $d_{\sigma_{+}}=d_{\sigma_{-}}=\sqrt{2}$. Thus, no Abelian sector exists on annulus $X^{\prime}$. If we consider an embedded annulus $X$ not surrounding any defect, then the information convex set $\Sigma(X)$ is a simplex with four extreme points. These extreme points correspond to Abelian sectors $1$, $e$, $m$, and $f$. The model is solvable, and explicit computation [^13] Similar computation can be done in some non-Abelian models, such as quantum double using “minimum diagram” technique [16, 36]. can be done to verify these statements. Putting aside the details, examples like this are nontrivial contexts for question (27).

### v4

@@ 2.4 Superselection sectors and fusion spaces
Information convex sets detect superselection sectors and fusion spaces. In 2D, an embedded annulus $X$ detects the anyon types ${\mathcal{C}}=\{1,a,b,\cdots\}$, and here $1$ denotes the vacuum sector. The 2-hole disk (Fig. 4 (a)) detects the fusion spaces. These follow from the structure theorems: the simplex theorem and the Hilbert space theorem in the original paper [14]. The simplex theorem states that the information convex set of the embedded annulus, $\Sigma(X)$, is a simplex:
@@
More generally, the simplex theorem works for any sectorizable region [15]. Sectorizable is a topological condition that says $S$ contains disjoint subsets $S_{1}$ and $S_{2}$, each of which can deform back to $S$ by a sequence of elementary steps of extensions. (The definition of “elementary step” is reviewed at the beginning of Section 2.3.) In 2D, immersed annuli and disks are the only sectorizable regions. A Möbius strip embedded in $RP^{2}$ is not sectorizable, even though it is the thickening of some loop.
@@ 2.5 Quantum dimensions
Let $Z$ be an immersed annulus. Let $Z=BCD$, where $BD=\partial Z$ is the thickened boundary and $C=Z\setminus\partial Z$ is the interior; each of $B$ and $D$ is a union of two disks, as illustrated in Fig. 5. For any extreme point $\rho^{h}_{Z}\in\Sigma(Z)$,
@@
Figure 5: The partition of an immersed annulus $Z=BCD$ used in defining quantum dimension (Definition 4). We first map the immersed annulus to a topological annulus by a homeomorphism indicated by the arrow. We then partition the topological annulus as illustrated.
@@
According to Lemma 7, any one of the conditions 1, 2, 3, and 4 can be used to determine if a state in $\Sigma(Z)$ is an extreme point associated with an Abelian superselection sector.
@@
First, suppose that $\rho_{Z}$ is an extreme point of $\Sigma(Z)$. According to the definition of quantum dimension, $\Delta(B,C,D)_{\rho}=0$ for the partition in Fig. 5. It follows from Lemma 15 in Appendix A that any one of conditions 1, 2, 3, 4 implies the rest of them. (Lemma 15 in the appendix is a more general statement about the equivalence of definitions for quantum dimension, works beyond 2D setups.) Second, none of 1, 2, 3, 4 is true if $\rho_{Z}=\sum_{h}p_{h}\,\rho^{h}_{Z}$ is a non-extreme point. This is because a non-extreme point has contribution $\sum_{h}p_{h}\ln\left(\frac{1}{p_{h}}\right)>0$ added to $\Delta$s and $I$ in conditions 1, 2, 3, 4. This contribution is always positive for non-extreme points. From these two considerations, we derive Lemma 7. ∎
@@ 3 Fun with figure-8
We emphasize that the conjecture is made in the context where the reference state is defined on the sphere $\mathbf{S}^{2}$. The axioms A0 and A1 are satisfied everywhere on the sphere. An alternative version of the conjecture requires the reference state $\sigma$ defined on an open disk. These two versions are equivalent due to the ability to “complete” the disk into a sphere, as commented earlier. Note, however, that the conjectured statement is false if we remove the requirement of A1 centered at two or more points on the sphere. The counterexample is a state with a pair of topological defects. If topological defects (e.g., that permute $e$ and $m$ in the toric code model, reviewed in Section 5.1) are present, then an annulus (either embedded or a figure-8) that winds the defect point around must have no Abelian sectors. We shall prove Conjecture 8 for Abelian anyon theory in Section 4.
@@
The weaker statement $|{\mathcal{C}}_{8}|=|{\mathcal{C}}|$, i.e., the number of extreme points of $\Sigma(X_{8})$ equals the number of anyon species, is true; we will discuss this in Section 5. If the two information convex sets $\Sigma(X_{8})$ and $\Sigma(X)$ are isomorphic in the most natural manner, we expect an invertible map between them that preserves the quantum dimensions. Then, in addition to $|{\mathcal{C}}|=|{\mathcal{C}}_{8}|$, we should have an Abelian sector of ${\mathcal{C}}_{8}$ that is mapped to the vacuum $1\in{\mathcal{C}}$. This is one way to motivate Conjecture 8. (It is a further question whether there is a “canonical vacuum” on ${\mathcal{C}}_{8}$.)
@@ 3.1 Tunneling trick
We introduce a tunneling trick. The essential observation is that tunneling a figure-8 through a strip can add a twist to the strip. If the “figure-8” region carries an Abelian sector, this process generates an isomorphism between information convex sets associated with homeomorphic topological regions in possibly distinct regular homotopy classes. We first explain the tunneling process at a topological level. After that, we discuss the details of entanglement bootstrap and explain what tunneling does to the quantum state (density matrices) and why an Abelian sector in ${\mathcal{C}}_{8}$ plays a role.
@@
What does the tunneling process in Fig. 7 mean in the entanglement bootstrap? It represents a transformation of the quantum state (density matrix) on the region containing the strip. Let the region containing the original strip be $A$ and the region containing the strip after tunneling be $A^{\prime}$, (see Fig. 8). Suppose, the state on “figure-8” $C$ is an extreme point $\rho^{\mu}_{C}$, with $\mu\in{\mathcal{C}}_{8}$. The transformation, as we shall discuss, must be a quantum channel that depends on $\mu$:
@@
Figure 8: Analyzing the attachment (detachment) of an immersed figure-8 annulus $C$ ($C^{\prime}$). Only part of $A$, $A^{\prime}$ is shown. (a) A partition for merging. (b) and (c) have special usage when the state on “figure-8” is an Abelian extreme point.
@@
The derivation of Eq. (22) is as follows. We first extend the state $\alpha_{ABC}\in\Sigma_{[\rho^{\mu}_{C}]}(ABC)$ to a state in $\Sigma(ABCD)$, as in Fig. 8(b). If $\mu$ is Abelian, by Lemma 7, we have $\Delta(B,C,D)_{\rho^{\mu}}=0$. Then by strong subadditivity, $I(A:C|B)_{\alpha}\leq\Delta(B,C,D)_{\rho^{\mu}}=0$. Because of its quantum Markov state structure, $\alpha_{ABC}$ can be recovered from $\alpha_{AB}$ by a quantum channel that is independent of $\alpha_{AB}$ (but depends on $\rho^{\mu}$). The fact that this $\alpha$-independent quantum channel reverses the partial trace ${\rm Tr}_{C}$ implies that $\Gamma^{\rm att}$ is an isomorphism. For the same reason, the detachment map $\Gamma^{\rm det}$ is an isomorphism too. (The state exists by merging. The reversibility follows from $I(A^{\prime}:C^{\prime}|B^{\prime})=0$.)
@@
Here, a 0-handle is a disk (the round part of the figure), and a 1-handle is the strip (thick lines in the figure). Both 0-handles and 1-handles are topological balls. The difference is the way they attach. This illustration will be convenient.
@@ 3.2 Abelian sector on “figure-8” implies 2D strong isomorphism
Components participating in the boundary connected sum (i.e., punctured tori and annuli) can be interchanged by sliding one across another; see the first deformation step [^9] This is an analog of a well-known trick in topology, called handle slides, which is useful to arrange handles into a certain order (canonical form), see e.g., Fig. 1 of [36]. of Fig. 9. Such a sliding trick works for both embedded regions and the more general immersed regions. After we arrange these components in the desired positions, we can deform them to 1-handles attached to the 0-handle, as shown in the second step in Fig. 9.
Figure 9: Step 1: We rearrange components in a boundary connected sum by sliding one component across another. The gray piece at the bottom is a large disk (0-handle). The way to slide the annulus (orange) from the right to the left across the punctured torus (blue) is indicated by the dashed line with an arrow. Step 2: We deform each component to 1-handles attached to the 0-handle. This technique works when these components are immersed.
It follows that we can smoothly deform $\Omega$ on $\mathbf{S}^{2}$ such that it looks like the left figure of Fig. 10. It is a 0-handle with $12k+m-$ 1 1-handles attached in the order (around the 0-handle) indicated in the figure. Some of the 1-handles have twists. Upon the removal of possible twists, $\Omega$ becomes $\Omega^{\star}$ in the right figure of Fig. 10.
@@
In $\Omega^{\star}$, the 1-handles form two groups: $(m-1)$ of them are separated from others (dark brown), and there are $k$ pairs of 1-handles (dark blue), the two 1-handles in each pair “cross” each other. Importantly, in both figures in Fig. 10, we can arrange the 1-handles in the same order, circling around the 0-handle (the chunky gray disk), thanks to the ability to slide the components around each other, as explained in the previous paragraph.
@@ 4 Proofs in the context of 2D Abelian anyon theory
We consider a sequence of operations applying to a state on “figure-8” $X_{8}$ carrying an arbitrary $\mu\in{\mathcal{C}}_{8}$; see Fig. 11. We denote this extreme point of $\Sigma(X_{8})$ as $\rho^{\mu}_{X_{8}}$. We draw $X_{8}$ in such a way that part of the “figure-8” annulus is blown up. The region is like a “watch with a twisted belt”. We shall explain the steps in Fig. 11 and the fact that they imply a useful condition, Eq. (29), on the $\rho^{\mu}_{X_{8}}$. Below are the details.
@@
The key ingredient of the above proof is that transporting anyon along the figure-8 does not change its type; here, we assume that the figure-8 region is immersed in a ball or a sphere. Again, we emphasize that the crucial input is the anyon transportation property in the neighborhood of the immersed region, and it is not essential to assume the immersion is in a ball or sphere if we know this property. This observation will be useful later in Section 5.
@@ 5 Transportation experiment
We discuss a thought experiment, “transportation experiment,” which reveals the physical connection between immersed annuli with topological defects. The entanglement bootstrap setup in this section is relaxed compared with that in Sections 3 and 4. Let us consider 2D for simplicity. We allow the reference state to have “defect points” where the axioms may be violated. (Alternatively, we consider a reference state defined on a manifold $N$ with an interesting topology, e.g., a torus or a $k$-hole disk.) The setup is illustrated in Fig. 12.
@@
We want to transport an anyon along a closed loop ${\mathcal{L}}$, so that the anyon goes back to the original position in space. In entanglement bootstrap, this is done by transporting a “test annulus” $X$, which is a sufficiently small embedded annulus that detects the anyon type by the information convex set $\Sigma(X)$. Suppose we transport the test annulus $X$ along a closed loop ${\mathcal{L}}$, so that it goes back to its original position, as illustrated in Fig. 12. We can assign an automorphism of the information convex set determined by the loop ${\mathcal{L}}$: [^11] Accurately speaking, the automorphism is determined by a loop ${\mathcal{L}}$ with a chosen orientation. We may denote an oriented loop as $\vec{\mathcal{L}}$. However, we shall use ${\mathcal{L}}$ for simplicity. Quantities of interest below, such as the subset of anyons not permuted by the transportation, do not depend on the orientation. (For a discussion of automorphisms of information convex sets in broader setups, see Appendix B.)
@@
In general, the automorphism $\Phi({\mathcal{L}})$ can permute extreme points of $\Sigma(X)$, and thus, it permutes anyon types. In other words, for each ${\mathcal{L}}$, we have an automorphism of labels $\varphi_{\mathcal{L}}:{\mathcal{C}}\to{\mathcal{C}}$. The automorphism is invariant under smooth deformations of loop ${\mathcal{L}}$ that do not pass the defect points. Explicit examples come from topological defects, which we recall in Section 5.1. Throughout this section, we consider loops ${\mathcal{L}}$ whose thickening ($X_{\mathcal{L}}$) is an immersed annulus. [^12] This excludes some loops on nonorientable manifolds. For instance, the noncontractible loop on $RP^{2}$ can be thickened to an Möbius strip but not an annulus. We define
@@ 5.1 Topological defects: an example
A well-known example of a topological defect is the one that permutes $e$ and $m$ in the toric code model [37]. The anyon types of the toric code are ${\mathcal{C}}=\{1,e,m,f\}$. Their quantum dimensions are $d_{1}=d_{e}=d_{m}=d_{f}=1$. Topological defects in this model are non-Abelian ${\mathcal{C}}^{\prime}=\{\sigma_{+},\sigma_{-}\}$ with $d_{\sigma_{+}}=d_{\sigma_{-}}=\sqrt{2}$. Anyons $e$ and $m$ are permuted when they are transported around the defect (either $\sigma_{+}$ or $\sigma_{-}$). While defects are most familiar in the context where global symmetries are present [38], it is understood that the phenomenon of permuting anyons does not go away under arbitrary local perturbations, not preserving any symmetry.
@@
The presence of a defect affects information convex sets. This is illustrated in Fig. 13 for the toric code example. The annulus $X^{\prime}$ surrounds the defect, and its information convex set $\Sigma(X^{\prime})$ has two extreme points. The quantum dimension for each extreme point, defined according to Definition 4, agrees with $d_{\sigma_{+}}=d_{\sigma_{-}}=\sqrt{2}$. Thus, no Abelian sector exists on the annulus $X^{\prime}$. If we consider an embedded annulus $X$ not surrounding any defect, then the information convex set $\Sigma(X)$ is a simplex with four extreme points. These extreme points correspond to Abelian sectors $1$, $e$, $m$, and $f$. The model is solvable, and explicit computation [^13] Similar computation can be done in some non-Abelian models, such as quantum double using the “minimum diagram” technique [16, 39]. can be done to verify these statements. Putting aside the details, examples like this are nontrivial contexts for question (33).